8-pack control system working group Meeting Notes

28 October 2003


In attendance: E. Andrikopoulos, R. Chestnut, J. Frisch, R. Hall, K. Kim, J. Nelson, S. Smith

Note taker/typer: J. Nelson



Status Items:


      Mods for better SIS resolution – JF/SS/&c

            Plans are still being decided, but might include amps on the input signals and tweaking the SIS electronics, such that full scale is +/- 100 mV instead of +/-1V.

      Trickle Down of Sled and Phase Numbers

            Tested and A-Ok!  Works great!


            SLED/LD Databases/Panel Mods (incl plots)

            Janice reported that there are some purple statuses at the bottom of the SLED panel, and suspects a misplaced colon.

            Proc loop panel – need an on/off button and get rid of the 4 buttons across the top from the left and replace with just one button that is a summary display of the 8 interesting values on the ADC >db panels

            Dm2k plot with just one ampl and phase – Janice started this, but had CVS problems.  Will also need a panel with ~40 buttons to call this display for all the A&P WFs.  Ron thinks we can force the x-scale, which might also help.

                  (Ron will release Janice’s panel mods and Janice will ask for more unix quota, hoping that’s the cause of her problem.)


      Hardware Deadman

            Steve will test the hardware hook to RF chassis.


      IQAs and PHITs

            Janice reported that 5 of 6 IQAs are out there, but IQA5 has a bad daughter board.  Also, when two IQA channels are stolen for the two load forward powers, we won’t bother to change the software.



            No change- still panel work and testing to do.  Cassel will be modifying the PLC in the next couple weeks and will most likely add status bit that we’ll want in EPICS.


      AB Purple Status

            (no change) The EPICS timeout was changed a week or so ago.  Judy will check cmlog to see if the frequency of the purple status has changed.


      Peak Power Meters

            Labview exists to pull out the waveforms, but has yet to stuff them into the CA server.  Maybe soon?


      Knobs for Controlling SLED Movers

            Janice made these and used them and they worked great.



Things to Test (tentative start date is 11/6)

      Calibrate IQAs and generator (get SIS testing for free)

            Calculate ADC conversion values

      Processing loop with real values

      RE/vac/proc loop/etc trips, also diode SW reset

      Check and test that all interesting channels are being archived appropriately.




      MacNair Digitizer Support

            Kukhee is almost done.  Maybe later this week we can start testing.


      Janice reported that the FONT SISes are working great.  Thanks!


      Joe has a request for all the space DAC channels in TR01.  No need for panels, just need database/PVs.  Ron will look into how many spares and setting up the database.


      Bob reported that both margetdata and taoutdata work ok for getting waveforms from Oracle, provided the datasets are sufficiently small.  Janice will give Bob the start date and PV names so he can get the data directly from Oracle and put it in a file for Janice.